# Formatting Instructions For NeurIPS 2024

### v1

@@ Appendix A The linearized GP of GCNs
$\Delta_{\alpha\beta}^{(l+1)}=$ $\sum_{\gamma\delta}H_{\alpha\beta,\gamma\delta}\Delta_{\gamma\delta}^{(l)}+\mathcal{O}((\Delta^{(l)})^{2})$ (24)
@@
$\lambda_{i}V_{\alpha\beta}^{(i)}=$ $\sum_{\gamma\delta}H_{\alpha\beta,\gamma\delta}V_{\gamma\delta}^{(i)}\,.$ (25)
@@
$\sum_{\alpha\beta}U_{\alpha\beta}^{(i)}V_{\alpha\beta}^{(j)}=$ $\delta_{ij}$ (26)
@@
$\Delta_{i}^{(l)}=$ $\sum_{\alpha\beta}U_{\alpha\beta}^{(i)}\Delta_{\alpha\beta}^{(l)}$ (27)
@@
Appendix B Analytical solution for expectation values
To evaluate our theory, we need to compute expectation values given in the form of 8 which we restate here for readability
where $h_{\gamma}^{(l)}$ and $h_{\delta}^{(l)}$ are zero mean random Gaussian variables with $\langle h_{\gamma}^{(l)}h_{\delta}^{(l)}\rangle=K_{\gamma\delta}^{(l)}$. This can be evaluated numerically for general nonlinearities $\phi(x)$. For simplicity, however, we choose $\phi(x)=\mathrm{erf}(\frac{\sqrt{\pi}}{2}x)$ in our experiments where the scaling factor in the $\mathrm{erf}$ is chosen such that $\frac{\partial\phi}{\partial x}(0)=1$. In this case, the expectation value can be evaluated analytically to be
as shown in [39].
@@ Appendix C Investigation of the complete graph model
where we used the known solution for the expectation value $C_{\mathrm{c}}$ for $\phi(x)=\mathrm{erf}(\sqrt{\pi}x/2)$ from Appendix B. Plugging this into Equation (43) lets us calculate $\frac{\partial c^{(l+1)}}{\partial c^{(l)}}\Big{|}_{c^{(l)}=1}$ and thus determine the transition to the non-oversmoothing phase. This is plotted as a red line in Figure 1 panel a) and b).
Appendix D Details of numerical experiments
To conduct our experiments we use NumPy [12], SciPy [37] (both available under a BSD-3-Clause License) and Scikit-learn (sklearn) [28] (available under a New BSD License). Computations were performed on CPUs, the requirements for the experiments are (approximately):
@@ Appendix D Details of numerical experiments
  • Figure 3: 2h on a single core laptop. In panel d), the last layer of finite-size GCNs is trained with the standard settings from sklearn.linear_model.SGDRegressor()
Appendix E Restriction of weight matrices in related work
In [2] which is based on [27] the authors study GCNs, however with a different shift operator than we have in our work. In both of these works the authors state that the GCN exponentially looses expressive power if $s\overline{\lambda}<1$ with $\overline{\lambda}$ being the maximal singular value/eigenvalue of the shift operator and $s$ being the maximal singular value of all weight matrices. Again, the maximal singular value is limited (dependent on $\overline{\lambda}$) which in our approach limits $\sigma_{w}^{2}$.

### v2

@@ Appendix B The linearized GP of GCNs
$\Delta_{\alpha\beta}^{(l+1)}=$ $\sum_{\gamma,\delta}H_{\alpha\beta,\gamma\delta}\Delta_{\gamma\delta}^{(l)}+\mathcal{O}((\Delta^{(l)})^{2})$ (26)
@@
$\lambda_{i}V_{\alpha\beta}^{(i)}=$ $\sum_{\gamma,\delta}H_{\alpha\beta,\gamma\delta}V_{\gamma\delta}^{(i)}\,.$ (27)
@@
$\sum_{\alpha,\beta}U_{\alpha\beta}^{(i)}V_{\alpha\beta}^{(j)}=$ $\delta_{ij}$ (28)
@@
$\Delta_{i}^{(l)}=$ $\sum_{\alpha,\beta}U_{\alpha\beta}^{(i)}\Delta_{\alpha\beta}^{(l)}$ (29)
@@ Appendix C Investigation of the complete graph model
where we used the known solution for the expectation value $C_{\mathrm{c}}$ for $\phi(x)=\mathrm{erf}(\sqrt{\pi}x/2)$ from Appendix A. Plugging this into Equation (43) lets us calculate $\frac{\partial c^{(l+1)}}{\partial c^{(l)}}\Big{|}_{c^{(l)}=1}$ and thus determine the transition to the non-oversmoothing phase. This is plotted as a red line in Figure 1 panel a) and b).
Appendix D Restriction of weight matrices in related work
The authors of [2] and [30] also study GCNs,; however they consider a different shift operator than in this work. In both of [2] and [30] the authors find that their GCN models exponentially loose expressive power if $s\overline{\lambda}<1$, with $\overline{\lambda}$ being the maximal singular value/eigenvalue of the shift operator and $s$ being the maximal singular value of all weight matrices. Again, the maximal singular value is limited (dependent on $\overline{\lambda}$), which in our approach translates to a limit on $\sigma_{w}^{2}$. While the authors notice that their bounds do not hold for large singular values $s$, they do not observe a non-oversmoothing phase in their models.
Appendix E Non-oversmoothing GCNs with non-stochastic shift operators
Here we investigate how increasing the weight variance can mitigate oversmoothing also in the case of the original shift operator proposed by Kipf and Welling in [19] being
$A_{\text{KW}}$ $=(D^{\prime})^{-1/2}\mathcal{A^{\prime}}(D^{\prime})^{-1/2}$ (48)
with $\mathcal{A}^{\prime}=\mathbb{I}+\mathcal{A}$ and $D_{ij}^{\prime}=\delta_{ij}\sum_{k}\mathcal{A}_{ik}^{\prime}$. This shift operator $A_{\text{KW}}$ is not row-stochastic, therefore the state $K_{\alpha\beta}=k$ is not a fixed point for $k\neq 0$, as can be seen from Equation (19). Thus we need to differentiate between two kind of fixed points: Either, all features are zero, for which $K_{\alpha\beta}=0$ and $\mu(X)=0$, or some $K_{\alpha\beta}\neq 0$ in which case we have $\mu(X)>0$. Importantly, for this shift operator, there is no intermediate case for which $K_{\alpha\beta}=k$ with $k\neq 0$. Consequently there is no oversmoothing regime with respect to the measure $\mu(X)$ where node features are non-zero. For an in depth study of this case (48), the definition of another oversmoothing measure $\mu^{\prime}$ incorporating the different values of $\sum_{\beta}(A_{\text{KW}})_{\alpha\beta}$ for different $\alpha$ may be more appropriate. For our purposes, it will suffice to analyze the shift operator (48) with the measure $\mu(X)$ from Equation (14).
Figure 6 panel a) shows how $\mu(X)$ in equilibrium increases for larger weight variances $\sigma_{w}^{2}$ for the shift operator (48). For comparison, we also show the results obtained with our row-stochastic shift operator (6). Thus we find that also for the shift operator $A_{\text{KW}}$ the pairwise distance between features can be increased by increasing the weight variance $\sigma_{w}^{2}$. The non-existence of an oversmoothed fixed point except for the special case $K_{\alpha\beta}=0$ which we argued for above is observed in panel b) and c): The oversmoothing measure $\mu(X)$ is zero for $A_{\text{KW}}$ if and only if all entries of the covariance matrix are zero, implying that all preactivations and thus also all features are zero. This is a qualitative difference to our shift operator $A$ for which we find equilibrium states with non-zero $\max_{\alpha}K_{\alpha\alpha}^{\text{eq}}$ but still $\mu(X)=0$.
Figure 6: Oversmoothing in GCN GPs with the commonly used shift operator (48) (called Kipf & Welling in the label) and our row-stochastic shift operator (6). a) Feature distance $\mu(X)$ in equilibrium dependent on weight variance $\sigma_{w}^{2}$ obtained from simulating the GCN GP priors for $L^{\text{eq}}=4,000$ layers. Shown are the averages (solid lines) and standard deviations (shaded areas) over $20$ CSBM initializations with $\lambda=1$, $d=5$ and $N=30$ nodes.b) Scatter plot of maximal variance of all nodes $\max_{\alpha}\{K_{\alpha\alpha}^{\text{eq}}\}$ and feature distance $\mu(X)$ in equilibrium for the same data as in plot a). c) Same as b) but zoomed into lower left corner.
Appendix F Details of numerical experiments
To conduct our experiments we use NumPy [13], SciPy [42] (both available under a BSD-3-Clause License) and Scikit-learn (sklearn) [31] (available under a New BSD License). The code is publicly available under https://github.com/bepping/non-oversmoothing-gcns. For our experiments with the Cora dataset, we use the readin methods from [19] which are available under a MIT license (Copyright (c) 2016 Thomas Kipf). Computations were performed on CPUs.Requirements for the experiments with synthetic data are (approximately):
@@ Appendix F Details of numerical experiments
  • Figure 3: 2h on a single core laptop. In panel d), the last layer of finite-size GCNs is trained with the standard settings from sklearn.linear_model.SGDRegressor().
  • Figure 5: Byproduct of Figure 2.
  • Figure 6: 10min on a single core laptop.
We also experimented on the real world benchmark dataset Cora [38]. This is a citation network with $2708$ nodes, representing publications, and $5429$ edges, representing the citations between them (we use undirected edges). The publications are divided into seven classes, and the task is to predict these classes for the unlabeled nodes. Node features are of 00/$1$ valued vectors indicating the absence/presence of words from a dictionary in the titles of the respective publication. The dictionary consists of $1433$ unique words. Requirements for the experiments with the Cora dataset are:
  • Figure 4: 1h on a single core laptop.
  • Figure 7: 15h on a single core laptop.
Appendix G Non-oversmoothing transition in the Cora dataset
For Figure 4 we numerically determined the transition to the non-oversmoothing regime for the Cora dataset. The transition was estimated by measuring the distance $\mu(\boldsymbol{X})$ at equilibrium (i.e., after many layers), and determining at which value of $\sigma_{w}^{2}$ it becomes larger than a small distance $\epsilon=10^{-5}$. The results of this experiment are shown in Figure 7. We find the critical point to be $\sigma_{w,\text{crit}}^{2}=1$ while using a step size of $\delta\sigma_{w}^{2}=0.01$.
Appendix H Transitioning between regimes during training
Figure 7: Node distance measure $\mu(\boldsymbol{X})$ at equilibrium obtained from computing the GCN GP prior for $L=4,000$ layers as a function of $\sigma_{w}^{2}$. The transition to the non-oversmoothing regime is estimated by checking where the node distance measure is larger than $\epsilon=10^{-5}$, marked as the red line.
In this section argue why we think it is possible to transition from the oversmoothing to the non-oversmoothing regime or vice versa during training.
In the GP limit of infinitely many hidden features with a finite amount of training data, the variance of weights of a neural network is the same before and after training. This is because weights only change marginally in this limit, also known as the lazy training regime [16, 10]. We will use this fact to argue that Langevin training is capable of transitioning from one regime to the other.
Langevin training is a gradient based training scheme with external noise and a decay term. The gradient flow equation is given as
$\frac{\mathrm{d}W_{ij}}{\mathrm{d}t}=-\gamma W_{ij}-\nabla_{W_{ij}}L+B_{ij}\,,$
where $W_{ij}$ are the weights to be trained, $L$ is the loss function and $B_{ij}$ denotes external white noise $\langle B_{ij}(t)B_{kl}(s)\rangle=\delta_{i,k}\delta_{j,l}\delta(t-s)D$ with strength $D$ where $\delta_{a,b}$ and $\delta(a-b)$ denote the Kronecker or Dirac delta, respectively. Both $\gamma$ and $D$ are parameters of this training scheme. It is known that the distribution of weights converges to the posterior weight distribution of a neural network GP with weight variance $\sigma_{w}^{2}=\frac{D}{2\gamma}$ in the infinite feature limit [25] which, as we have noticed above, is the same as the prior distribution.
Since the Langevin distribution converges to the GP weight posterior for any initial distribution, one can imagine an initial distribution with a variance that initializes the network in the oversmoothing regime, while the parameters $\gamma$ and $D$ are chosen such that $\sigma_{w}^{2}=\frac{D}{2\gamma}>\sigma_{w,\text{crit}}^{2}$ implying that after training most weight realizations will be in the non-oversmoothing regime. Thus, in this case the GCN would have transitioned from one regime to the other. While this argument is made in the limit of infinitely many hidden features, we think that qualitatively similar results are possible for a large but finite number of hidden feature dimensions. The transition in the reverse direction is possible by the same argument only with the initial and the final variances exchanged.